GALOIS REPRESENTATIONS AND MODULAR FORMS

Kenneth A. Ribet

Abstract. In this article, I discuss material which is related to the recent proof of Fermat’s Last Theorem: elliptic curves, modular forms, Galois representations and their deformations, Frey’s construction, and the conjectures of Serre and of Taniyama–Shimura.

Typeset by AMS-TEX